Calculate Δ?, Δ?, Δ????, ??? Δ????? when 4.31 moles of H2O(g) condenses at 58 oC at 1.0 atm until it is a liquid. Assume that all heat capacities and enthalpy changes are constant with values of: ??,? = 75.3 ? ? ??? , ??,? = 35.9 ? ? ??? , ???? = 100 ?, Δ???? 40.7 ?? ??�
Answer #1

Reaction for the condensation: H2O(g) <----> H2O(l)

ΔCp = (Cp)H2O(l) - (Cp)H2O(g) = 75.3 - 35.9 = 39.4 J/K.mol

The change in temperature (ΔT) = 100 - 58 = 42 oC or 42 K

The no. of moles of H2O (n) = 4.31 mol

Now, ΔH = n*ΔCp*ΔT = 4.31 mol * 39.4 J/K.mol * 42 K = 7132.188 J or 7.132 kJ

Now, ΔS = ΔH/T = 7132.188 J/(58+273) K = 21.547 J/K